Problem 4 Let X be a discrete random variable with probability mass function fx(x), and let t be a function. Define Y = t(X): that is, Y is the randon variable obtained by applying the function t to the value of X Transforming a random variable in this way is frequently done in statistics. In what follows, let R(X) denote the possible values of X and let R(Y) denote the possible values of To compute E[Y], we could irst find its probability mass function fy(v), and then compute ER(Y) But the Rule of the Lazy Statistician tells us that it is sufficient to compute Σ t(z) . Ix(z). TER(X) We will prove that this is the case a) Let () t)y. Write a formula for fy () using t b) Using your formula for fy (), prove that JER(Y) rER(X)
Answer #1

Let us begin with the discrete random variable X with probability mass function  fx() and here as given

Y=toX=t(X)

where the function is such that

t : mathbb{R} ightarrow mathbb{R}

a) 1st part,

t-1 (y) = {x : t(x) = y}

The probability mass function f_{Y(y)} for  Y = t(X)

can be written in the form given below

fy(y) = ret-(y)

b) 2nd part,

fy(y) = ret-(y)

sum y.f_{Y}(y)=sum t(x).f_{X}(x)

where the range will change from ret-(y) to reR(X) and we know t(x) = y

sum y.f_{Y}(y)=sum t(x).f_{X}(x) along with the range reR(X)